# An overview of meta-analytic methods for economic research

### v1

@@ Abstract
Meta-analysis is the use of statistical methods to combine the results of individual studies to estimate the overall effect size for a specific outcome of interest. The direction and magnitude of this estimated effect, along with its confidence interval, provide insights into the phenomenon or relationship being investigated. As an extension of the standard meta-analysis, meta-regression analysis incorporates multiple moderators representing identifiable study characteristics into the meta-analysis model, thereby explaining some of the heterogeneity in true effect sizes across studies. This form of meta-analysis is especially designed to quantitatively synthesize empirical evidence in economics. This study provides an overview of the meta-analytic procedures tailored for economic research. By addressing key challenges, including between-study heterogeneity, publication bias, and effect size dependence, it aims to equip researchers with the tools and insights needed to conduct rigorous and informative meta-analytic studies in economics and related disciplines.
@@ 6 Effect Size Dependence
Ignoring or misspecifying dependence in meta-analytic data can lead to invalid estimates of the mean effect size (or the meta-regression coefficients) and their standard errors, resulting in inflated Type I error rates [Becker, 2000, Olkin & Gleser, 2009]. Several strategies have been employed to address statistically dependent effect sizes. The simplest of these involve ignoring the dependence among effect size estimates, selecting only one effect size estimate per study, computing a single summary effect size per study by averaging the effect size estimates clustered within studies, and shifting the unit of analysis. Although these strategies can effectively remove the dependence among effect size estimates, they result in a loss of information (see Rosenthal & Rubin, 1986, Cooper, 1998, Marín-Martínez & Sánchez-Meca, 1999, Becker, 2000, Jackson et al., 2011, Scammacca et al., 2014, Cheung, 2014, 2015a, Borenstein et al., 2021). A more principled approach is to explicitly model the correlations among effect size estimates using multivariate meta-analytic methods. The challenge with this approach is that it generally requires knowledge of the covariance structure of within-study effect size estimates, which is rarely reported in primary studies (see Raudenbush et al., 1988, Kalaian & Raudenbush, 1996, Becker, 2000, Olkin & Gleser, 2009, Jackson et al., 2011, Mavridis & Salanti, 2013). Fortunately, two alternative approaches—the robust variance estimation (RVE) and the three-level meta-analysis model—circumvent this issue. These recently developed approaches are discussed in the following subsections.
6.1 Robust Variance Estimation (RVE)
@@ 7 Concluding Remarks
Meta-analysis, a specific type of systematic review, is a robust statistical method used to combine quantitative findings from multiple studies, providing a comprehensive estimate of the overall or average effect size for a specific outcome of interest. The direction and magnitude of the estimated effect, accompanied by its confidence interval and hypothesis test results, provide insights into the phenomenon or relationship being investigated.
As a pivotal tool for synthesizing empirical evidence across various scientific disciplines, including economics, meta-analysis addresses the challenges posed by the exponential growth of research output. Individual studies often yield conflicting findings due to variations in data sources, methodologies, and contextual factors. By systematically integrating these results, meta-analysis bridges such gaps, offering a more robust and generalized understanding of economic phenomena. Its structured approach enhances statistical power, refines effect size estimates, and supports evidence-based decision-making.
A key strength of meta-analysis lies in its ability to address between-study heterogeneity—the differences in true effect sizes across studies. By employing meta-regression analysis, researchers can incorporate moderators representing study characteristics, thereby identifying the underlying sources of variability. This nuanced approach allows economists to draw more precise conclusions about the factors influencing economic relationships. For instance, it enables the examination of how varying labor market conditions shape the effectiveness of training programs or the differential impacts of policy interventions across regions and contexts.
@@
Another critical aspect of meta-analysis is the treatment of effect size dependence. Effect size dependence refers to the statistical interdependence of effect sizes reported across or within studies, which violates the assumption of independence that underlies traditional meta-analytic methods. Failing to account for this dependence can lead to biased inferences. In economic datasets, multiple effect sizes are often reported within a single study, resulting in non-independence of observations. Advanced methods, such as multivariate meta-analysis and robust variance estimation, address this complexity by adjusting for the correlation between effect sizes, ensuring that the synthesized results provide a more accurate reflection of the true relationships under investigation.
This study reviews the key tools and techniques for applying meta-analysis in economics, with particular emphasis on addressing fundamental challenges such as publication bias, between-study heterogeneity, and effect size dependence. By exploring these methods, the study offers a valuable resource for researchers seeking to conduct robust meta-analyses in economics. It also demonstrates how advanced techniques, such as meta-regression, can yield actionable policy insights by uncovering the underlying factors contributing to variability in economic research. Ultimately, this work advances the methodological development of meta-analysis in economics, equipping researchers with the necessary knowledge to effectively apply meta-analytic methods, thereby enhancing the reliability of empirical findings and supporting evidence-based policy formulation.

### v2

@@ Abstract
Meta-analysis employs statistical techniques to synthesize the results of individual studies, providing an estimate of the overall effect size for a specific outcome of interest. The direction and magnitude of this estimate, along with its confidence interval, offer valuable insights into the underlying phenomenon or relationship. As an extension of standard meta-analysis, meta-regression analysis incorporates multiple moderators—capturing key study characteristics—into the model to explain heterogeneity in true effect sizes across studies. This study provides a comprehensive overview of meta-analytic procedures tailored to economic research, addressing key challenges such as between-study heterogeneity, publication bias, and effect size dependence. It equips researchers with essential tools and insights to conduct rigorous and informative meta-analyses in economics and related fields.
@@ 4 Publication Bias
A variety of methods methods and techniques have been developed for the identification of and correction for publication bias in meta-analyses. [^10] For a detailed review of methods for detecting, assessing, and correcting for publication bias, see Vevea et al. [2019]. Some of the most common approaches are introduced in the following. Stanley et al. [2010] proposed the “Top 10” approach, where only the most precise 10% of reported estimates are used in meta-analyses. This approach mitigates publication bias by emphasizing high-precision studies, which are less prone to selection bias. By focusing on precision, the “Top 10” approach reduces the influence of less reliable studies, thereby enhancing the accuracy and reliability of meta-analytic conclusions. Stanley et al. [2017] proposed a meta-analytic method designed to mitigate selective reporting bias by focusing on estimates with adequate statistical power, typically defined as having at least 80% power to detect a nonzero effect. The Weighted Average of Adequately Powered (WAAP) method computes a weighted average of effect sizes using optimal weights based on the inverse of variance but applies these weights exclusively to estimates that meet the power threshold. By focusing on adequately powered studies, WAAP diminishes the influence of smaller, less reliable studies that may exaggerate effect sizes due to selective reporting. Overall, WAAP generally outperforms traditional meta-analytic models by significantly reducing selective reporting bias while maintaining strong statistical properties when no bias is present. Andrews & Kasy [2019] proposed two approaches to identify the probability of publication conditional on a study’s results. The first uses data from systematic replications of original studies, while the second uses data from meta-studies. The authors initially propose methods for calculating bias-corrected estimators and confidence sets for the parameters of interest when the conditional probability of publication is known. They then demonstrate how this conditional publication probability can be nonparametrically identified using data from replication studies and meta-analyses. Bom & Rachinger [2019] introduced the endogenous kink (EK) meta-regression model as a novel approach for correcting publication bias in meta-analysis. The EK model fits a piecewise linear regression of effect size estimates on their standard errors, featuring a kink at an endogenous cutoff value of the standard error. This cutoff represents the point below which publication selection is unlikely, as studies are deemed sufficiently significant, and increases in standard errors do not influence publication likelihood. Above this cutoff, publication selection becomes more likely, resulting in a positive linear relationship. The model endogenously determines this cutoff based on a first-stage estimate of the true effect size and a predefined threshold of statistical significance. Stanley et al. [2021] proposed three tests for detecting publication selection bias based on excess statistical significance (ESS): the Proportion of Statistical Significance Test (PSST), the Test of Excess Statistical Significance (TESS), and their combination (TESSPSST). These tests explicitly account for between-study heterogeneity by incorporating it into the formulas for expected and excess statistical significance calculations. The PSST assesses whether the observed proportion of statistically significant findings is notably larger than what could theoretically be expected given the mean statistical power. If it is, this is taken as an indication of publication bias. The TESS evaluates whether the proportion of ESS exceeds 5%, a threshold commonly accepted as the rate of false positives (i.e., Type I errors). The combined TESSPSST test identifies publication selection bias if either TESS or PSST is significant, thereby enhancing detection power beyond conventional methods such as the Egger test.
@@
McShane et al. [2016] reviewed model selection methods and conducted an extensive simulation study to assess different aspects of their performance, such as estimation accuracy and confidence interval coverage. The results indicate that while the p-curve and p-uniform methods perform reasonably well in restrictive settings (characterized by rigid publication/selection rules and homogeneous effect sizes), they do not perform as well as the original Hedges [1984] approach, due to the alternative estimation strategies they employ. Moreover, these approaches perform poorly in more realistic settings (involving more flexible publication rules and heterogeneous effect sizes), whereas variants of the Hedges method exhibit better performance. The authors conclude that, given the idealized assumptions underlying selection methods and the sensitivity of effect size estimates to these assumptions, such methods should not be used solely to obtain a single adjusted effect size estimate, but rather as part of a broader sensitivity analysis to explore the impact of different forms and degrees of publication bias.
As has been discussed so far, there is a wide range of methods available for detecting and correcting publication bias, each with its own strengths and limitations. However, in practice, researchers often lack a clear understanding of the data-generating process and do not have sufficient information to confidently select from these methods. Furthermore, this broad array of methods can often lead to conflicting conclusions. The combination of uncertainty regarding the data-generating process and these contradictions can create a “breeding ground” for confirmation bias, where researchers may inadvertently choose methods that align with their desired outcomes. This flexibility in method selection can significantly inflate the rate of false positives, posing a serious challenge to conventional meta-analytic methods [Maier et al., 2023, Bartoš et al., 2023].
Given these challenges, one promising alternative is to explicitly combine different models and allow the data to determine the weight of each model based on its predictive accuracy for the observed data. In this context, Maier et al. [2023] introduced a robust Bayesian meta-analysis (RoBMA) that model-averages selection models, fixed effects, and random effects models to address publication bias. By averaging over 12 models, RoBMA accounts for the uncertainty inherent in selecting meta-analytical models and is more robust to model misspecification. It provides quantifiable evidence for the absence of publication bias, performs well even under high between-study heterogeneity, and reliably avoids false positives. RoBMA also allows for the sequential updating of evidence as new studies become available, effectively addressing concerns about accumulation bias. The method outperforms other approaches in simulations and real data applications. Although RoBMA tends to underestimate effect sizes when the alternative hypothesis is true and overestimate them in the presence of publication bias, these biases are slight and disappear with larger sample sizes, particularly when the majority of the weight is assigned to the correct models.
Bartoš et al. [2023] extended the RoBMA framework by integrating PET-PEESE regression models (as well as one-sided weight functions) with p-value-based selection models, leading to the development of RoBMA-PSMA (publication selection model-averaging). This extension optimally combines both approaches to correct for publication bias, resolving the tension between them and enabling inference through a model-averaged weighting scheme that reflects how well each model fits the data. Bartoš et al. [2023] demonstrated that RoBMA-PSMA outperforms previous methods in meta-analyses for which a gold standard is available and performs well in simulations. However, the method exhibits limitations in settings with prominent p-hacking, where it tends to overestimate effect sizes. Overall, Bayesian model averaging, in the form of RoBMA-PSMA, significantly improves both PET-PEESE and selection models by reducing the MSE and bias of PET-PEESE, as well as the MSE of the selection models. Importantly, RoBMA-PSMA accounts for uncertainty regarding the type of publication bias and seamlessly integrates the strengths of both approaches. Moreover, the framework’s flexibility allows researchers to specify alternative prior distributions or incorporate additional models based on prior knowledge. Both RoBMA and RoBMA-PSMA can be implemented using the latest version of the ‘RoBMA’ R package.
@@ 6 Effect Size Dependence
Ignoring or misspecifying dependence in meta-analytic data can lead to invalid estimates of the mean effect size (or the meta-regression coefficients) and their standard errors, resulting in inflated Type I error rates [Becker, 2000, Olkin & Gleser, 2009]. Several strategies have been employed to address statistically dependent effect sizes. The simplest of these involve ignoring the dependence among effect size estimates, selecting only one effect size estimate per study, computing a single summary effect size per study by averaging the effect size estimates clustered within studies, and shifting the unit of analysis. Although these strategies can effectively remove the dependence among effect size estimates, they result in a loss of information (see Rosenthal & Rubin, 1986, Cooper, 1998, Marín-Martínez & Sánchez-Meca, 1999, Becker, 2000, Jackson et al., 2011, Scammacca et al., 2014, Cheung, 2014, 2015a, Borenstein et al., 2021). A more principled approach is to explicitly model the correlations among effect size estimates using multivariate meta-analytic methods. The challenge with this approach is that it generally requires knowledge of the covariance structure of within-study effect size estimates, which is rarely reported in primary studies (see Raudenbush et al., 1988, Kalaian & Raudenbush, 1996, Becker, 2000, Olkin & Gleser, 2009, Jackson et al., 2011, Mavridis & Salanti, 2013). Fortunately, three common approaches—the robust variance estimation (RVE) approach, the multilevel meta-analysis model, and the generalized weights (GW) approach—address this issue. These recently developed methods are discussed in the following subsections.
6.1 The Robust Variance Estimation (RVE)
@@ 6.2 The Multilevel Model
6.3 The Generalized–Weights (GW) Approach
Meta-studies are often conducted on empirical outcomes derived from studies that use overlapping samples. This is particularly common in fields such as economics and finance, which rely heavily on aggregated observational data. In these fields, the same or similar datasets are reused across multiple studies. This characteristic of meta-analysis data is referred to as sample overlap [Bom & Rachinger, 2020, 2024]. Bom & Rachinger [2020] demonstrate that sample overlap creates a dependence structure among primary estimates, which can result in asymptotic biases, significant efficiency losses, and increased rates of false positives at the meta-analysis level. To address this issue, they proposed the generalized-weights (GW) meta-estimator, which follows a two-step procedure. The first step involves constructing the complete variance-covariance matrix to capture the correlation structure between primary outcomes. This matrix can be feasibly estimated using readily available information from primary studies, such as sample sizes, standard errors, and the number of overlapping observations. In the second step, the estimated matrix is used to optimally weight each primary outcome based on its contribution of independent (nonoverlapping) sampling information. Bom & Rachinger [2020] employ Monte Carlo simulations to quantify the efficiency gains of the GW meta-estimator relative to conventional meta-estimators and to demonstrate how the GW meta-estimator brings the false positive rate closer to its nominal level. The GW meta-estimator is a simple and versatile tool designed to handle all types of sample overlap, whether within a study or across studies. It can accommodate complex cases without the need for additional estimation, provided that the necessary information to approximate the variance-covariance matrix of the estimates can be extracted from the primary studies, which is typically available in most applications.
The GW estimator provides several significant advantages over cluster-robust methods and multilevel models in meta-analysis, particularly when dealing with sample overlap. Unlike cluster-robust methods, which require defining disjoint clusters of correlated (overlapping) estimates and assuming a constant within-cluster correlation, the GW estimator addresses these challenges by utilizing exogenously determined overlap information. This method provides a more flexible way to correct for overlap-induced dependencies, especially when clusters are small, unbalanced, or irregular in size–situations in which cluster-robust methods tend to perform poorly. While multilevel models enhance efficiency by capturing hierarchical dependencies, they are limited by the need to define a nested structure and estimate numerous parameters, which can be technically demanding and impractical in complex overlap scenarios. In contrast, the GW estimator simplifies the process by eliminating the need for joint parameter estimation. Furthermore, the GW method addresses both inference and efficiency concerns by effectively utilizing overlap information, making it especially useful for complex and unbalanced overlap structures. Although it does not directly address all types of dependency, it can complement other methods, including multilevel and cluster-robust models, in meta-analytic applications [Bom & Rachinger, 2020, 2024].
Sample overlap is common in economics meta-analysis and can give rise to different structures of estimate dependence, ranging from balanced to unbalanced, and simple to complex. If not properly addressed, sample overlap can result in efficiency losses and inflated false positive rates in meta-analyses. To tackle these issues, Bom & Rachinger [2024] demonstrate the application of the GW method in meta-analytic research within economics, addressing common practical challenges such as variations in data aggregation levels, estimation techniques, and effect size metrics. They derive explicit covariance formulas for various overlapping scenarios, assess the accuracy of the covariance approximations, and employ Monte Carlo simulations to demonstrate how the method improves efficiency and reduces false positive rates. The GW estimator is versatile and applicable to both simple meta-analysis models and meta-regression models. Bom & Rachinger [2024] have developed Stata and R codes that fully automate the implementation of the GW estimator in meta-analyses, particularly in cases involving sample overlap. These software codes, along with detailed instructions, are freely available for download at https://osf.io/g5t2j.
@@ 7 Concluding Remarks
Meta-analysis, a specific type of systematic review, is a robust statistical method used to combine quantitative findings from multiple studies, providing an estimate of the overall or average effect size for a specific outcome of interest. The direction and magnitude of the estimated effect, along with its confidence interval and hypothesis test results, provide valuable insights into the phenomenon or relationship under investigation.
As an essential tool for synthesizing empirical evidence across various scientific disciplines, including economics, meta-analysis plays a pivotal role in addressing the challenges posed by the rapid growth of research output. Individual studies often yield conflicting results due to variations in data sources, methodologies, and contextual factors. By systematically combining these results, meta-analysis bridges such gaps and offers a more precise and generalized understanding of economic phenomena. Its structured approach enhances statistical power, refines effect size estimates, and supports evidence-based decision-making.
A key strength of meta-analysis lies in its ability to address between-study heterogeneity—the differences in true effect sizes across studies. By employing meta-regression analysis, researchers can incorporate moderators representing study characteristics, thereby identifying the underlying sources of variability. This nuanced approach allows economists to draw more accurate conclusions about the factors influencing economic relationships. For example, it facilitates the investigation of how varying labor market conditions shape the effectiveness of training programs, or how policy interventions produce differential impacts across regions and contexts.
@@
Effect size dependence, another critical issue in meta-analysis, refers to the statistical interdependence of effect sizes reported across or within studies, which violates the assumption of independence in traditional meta-analytic methods. Failing to account for this dependence can lead to distorted conclusions. In economic research, effect size dependence often arises due to factors such as multiple effect sizes within a single study or between-study sample overlap. Advanced techniques, such as robust variance estimation (RVE), the multilevel model, and the generalized weights (GW) meta-estimator, account for the dependence among primary estimates, ensuring that the synthesized results more accurately reflect the true relationships under investigation.
This study provides a comprehensive review of key tools and techniques for applying meta-analysis in economics, with a particular emphasis on addressing critical challenges such as publication bias, between-study heterogeneity, and effect size dependence. It also discusses the strengths and weaknesses of various methods. In conclusion, the study equips researchers with the necessary knowledge to apply these methods effectively, thereby enhancing the reliability of empirical evidence and supporting evidence-based policy development.